OpenStudy (anonymous):

Find the derivative of f(x) = 8/x at x = -1

OpenStudy (radar):

Express it as: \[f(x) = 8x ^{-1}\]Does that help?

OpenStudy (radar):

Just use the "power" rule.

OpenStudy (radar):

No monolog for me!

OpenStudy (anonymous):

@radar what is the power rule

OpenStudy (radar):

Lets say that you have something like this: \[f(x) = ax ^{b}\]and you want to find the derivative. Note that the variable is to the "b" power. The power rule says that you get the derivative by multiplying the value of the power times the value of the coefficient getting the new coefficient. The new power will equal the old power minus 1 or this:\[x' =bax ^{(b-1)}\]

OpenStudy (radar):

in "your" problem, a = 8. amd b = -1 just plug in the values and simplify.

OpenStudy (radar):

I will standby, let me know what you get.

OpenStudy (anonymous):

-8?

OpenStudy (radar):

That is a good start -8 will be the start of a solution,.......but what power is the x now? Review my first post. The power of the original function is -1 \[1/x = x ^{-1}\]

OpenStudy (radar):

What is \[x ^{(-1-1)}\]

OpenStudy (radar):

You can now finish on your own. as now it is just math/arithmetic...good luck in your studies.

OpenStudy (radar):

Getting \[-8x ^{-2} =-8/x ^{2}\]

OpenStudy (radar):

Now for the final answer....substitute from the problem (x = -1) getting -8/-1^2) = -8/1 = -8